OpenStudy (anonymous):

The greenhouse club is purchasing seed for the lawn in the school courtyard. The club needs to determine how much to buy. Unfortunately, the club meets after school, and students are unable to find a custodian to unlock the door. Anthony suggests they just use his school map to calculate the area that will need to be covered in seed. He measures the rectangular area on the map and finds the length to be 10 inches and the width to be 6 inches. The map notes the scale of 1 inch representing 7 feet in the actual courtyard. What is the actual area in square feet? (Enter just the number, no units)

OpenStudy (anonymous):

@Vocaloid

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

The school map shows the rectangle to be 10 in. by 6 in. Also, the school map has a scale of 1 inch (map) = 7 ft (real). If 1 in. represents 7 ft, 10 inches represent how many feet?

OpenStudy (anonymous):

70

OpenStudy (mathstudent55):

Correct. Now for the width. If 1 in. (map) = 7 ft (real), then 6 in. represents what real length?

OpenStudy (anonymous):

42

OpenStudy (mathstudent55):

Map Real 1 in. = 7 ft 2 in. = 2 * 1 in. = 2 * 7 ft = 14 ft 3 in. = 3 * 1 in. = 3 * 7 ft = 21 ft etc. 6 in. = 6 * 1 in. = 6 * 7 ft = 42 ft etc. 10 in. = 10 * 1 in. = 10 * 7 ft = 70 ft

OpenStudy (mathstudent55):

Correct. That means the real length and width of the courtyard are 70 ft and 42 ft. Now you need to find the area of a rectangle with length 70 ft and width 42 ft. Do you know how to find the area of a rectangle?

OpenStudy (anonymous):

2,940

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Correct A = L * W = 70 ft * 42 ft = 2,940 ft^2 They want only the number, so just write 2940m, and leave out the units ft^2.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

so just write 2940

OpenStudy (mathstudent55):

You're welcome. Good job! You actually did all the work.

OpenStudy (mathstudent55):

Yes

OpenStudy (mathstudent55):

Notice the last line of the problem: (Enter just the number, no units)